# TIP-Editor: An Accurate 3D Editor Following Both Text-Prompts And Image-Prompts

### v2

@@ Abstract.
submissionid: 820copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 28–August 01, 2024; Denverisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
@@ 2. Related Works
Text-to-image (T2I) diffusion models  (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022a), trained on large-scale paired image-text datasets, have gained significant attention since they can generate diverse and high-quality images that match the complicated text prompt. Instead of directly generating images from scratch, another popular and closely related task is to edit the given image according to the text prompt (Meng et al., 2021; Couairon et al., 2022; Kawar et al., 2022; Hertz et al., 2022; Avrahami et al., 2022; Brooks et al., 2022).
Another popular task is object/concept personalization, which aims at generating images for a specified object/concept defined in the given image collection. Textual Inversion (TI) (Gal et al., 2022a) optimizes special text token(s) in the text embedding space to represent the specified concept. DreamBooth (Ruiz et al., 2022) fine-tunes the entire diffusion model with a class-specific prior preservation loss as regularization. In general, DreamBooth generates higher-quality images since it involves a larger amount of updated model parameters (i.e. the whole UNet model). However, all the aforementioned methods do not support generating images containing multiple personalized objects simultaneously.
@@
Recently, text-driven radiance field editing methods have gained more and more attention for their editing flexibility and accessibility. For example, Instruct-NeRF2NeRF (Haque et al., 2023) employs an image-based diffusion model (InstructPix2Pix (Brooks et al., 2022)) to modify the rendered image by the users’ instructions, and subsequently update the 3D radiance field with the modified image. DreamEditor (Zhuang et al., 2023), 3D Paintbrush (Decatur et al., 2023) and Vox-E (Sella et al., 2023) enable better local editing by adopting explicit 3D representations (i.e. mesh and voxel), where the editing region is automatically determined by the 2D cross-attention maps. GaussianEditor (Chen et al., 2023a; Fang et al., 2023) adopts GS as the scene representation and incorporates 3D semantic segmentation (Kirillov et al., 2023a; Cen et al., 2023) to facilitate efficient and precise scene editing. However, these text-driven approaches lack precise control over the specified appearance and position of the editing results.
@@ 4.1. Stepwise 2D Personalization
In general, our stepwise personalization of the pre-trained T2I model (i.e., Stable Diffusion (SD) (Rombach et al., 2022a)) is based on DreamBooth (Ruiz et al., 2022), but with two significant modifications. These changes are essential to personalize both the existing scene and the novel content in the reference image. First, in the 2D personalization of the existing scene, we propose an attention-based localization loss to enforce the interaction between the existing and the novel content specified by the provided 3D bounding box (e.g., sunglasses on the forehead, see Fig. 7). Note that the reference image is not involved in this step. Second, in the 2D personalization of the novel content, we introduce LoRA layers to better capture the unique characteristics of the specified item in the reference image.
@@ 5.3. Comparisons with State-of-the-Art Methods
5.4. Ablation Study
Ablation studies on the stepwise 2D personalization
@@ 6. Conclusion and Limitations
Appendix A More Experimental Results
Figure A.1. The texture details of the rendered image $I_{c}$ from the coarse 3DGS are further enhanced by the denoising process.
Denoising results. Fig. A.1 demonstrates the texture details of the rendered image $I_{c}$ from the coarse 3DGS and its corresponding enhanced image after the denoising process.
Figure A.2. Coarse editing results optimized with different $\gamma$. We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row). Specifically, only using global SDS loss $\mathcal{L}^{G}_{SDS}$ results in obvious artifacts in the editable region. Coarse editing results optimized with different $\gamma$. We demonstrate the rendering images of the edited Gaussian $\mathcal{G^{\mathcal{B}}}$ (top row) and all Gaussians in the scene $\mathcal{G}$ (bottom row). As the $\gamma$ value increases, global SDS loss $\mathcal{L}^{G}_{SDS}$ obviously results in more artifacts in the editable region. On the contrary, the decreasing of $\gamma$ results in fewer artifacts, but increases the risk of inaccurate object placement and an unnatural color discrepancy between the background and the novel content since the context information is missing during optimization. Generally, setting $\gamma=0.5$ effectively balances $\mathcal{L}^{G}_{SDS}$ and $\mathcal{L}^{L}_{SDS}$, reducing the noise Gaussian while correctly placing the sunglasses.
Influence of the weighting parameter in the coarse stage. Fig. A.2 demonstrates the coarse editing results optimized with different $\gamma$ in Eq. 8. As the $\gamma$ value increases (i.e. increasing the importance of the global SDS loss), more and more obvious artifacts appear in the editable region (i.g. the sunglasses). On the other hand, using too small $\gamma$ values results in fewer artifacts in the sunglasses. However, the final composited image becomes unnatural (e.g. the color gap between the original scene and the new object, the placement of the new object) due to lack of context information during optimization. Generally, setting $\gamma=0.5$ in Eq. 8 strikes a good balances, effectively reducing the noise in the editable Gaussians $\mathcal{G}^{\mathcal{B}}$, correctly placing the sunglasses, and producing harmony colors.
Appendix B Implementation Details
Figure B.1. The generation process of $I_{gt}$ in Sec. 4.3 of the main text. $I_{c}$ denotes the rendered image of the coarse 3DGS. Firstly, we add noise on $I_{c}$ to obtain $I_{c}^{noised}$ and then utilize the personalized T2I model $\epsilon_{\phi_{2}}$ as a denoising network and obtain $I_{c}^{d}$. Secondly, we obtain the binary instance mask $M^{inst}$ of the edited object/part by rendering only the editable Gaussians $\mathcal{G}^{\mathcal{B}}$ and thresholding its opacity mask. Finally, the pseudo GT image $I_{gt}$ is obtained as: $I_{gt}=M^{inst}\odot I_{c}^{d}+(1-M^{inst})\odot I_{bg}$
We provide more implementation details that cannot be included in the main paper due to space. All our experiments are conducted on a single NVIDIA Tesla A100 GPU.
B.1. Optimization of the initial GS scene
In our experiments, we employ 3D Gaussian Splatting (GS) (Kerbl et al., 2023) to represent the 3D scenes. Given images shot from different views of the target scene, we first use COLMAP (Schönberger and Frahm, 2016)) to estimate their corresponding camera parameters. Then, we train the initial GS scene with the default hyper-parameters in the official 3DGS repository (Snosixtyboo, [n. d.]).
B.2. Details of the stepwise 2D personalization strategy
We use the publicly available Stable Diffusion V2.1 (Rombach et al., 2022b) model provided on Hugging Face.
The scene personalization step. We sample scene images with the camera poses covering all elevation and azimuth angles with an interval of $30^{\circ}$ within a predefined range (varied according to the visual scope of the original scene images). We use BLIP-2 (Li et al., 2023) model to obtain the initial scene text prompt, and insert a special token $V_{1}$ in front of the noun to describe this specific scene, resulting in a scene-specific text prompt $y$ during fine-tuning. For the prior preservation loss, we randomly generate 200 images at resolution $512\times 512$ using the same initial scene text $y^{*}$ as input. We use the AdamW optimizer with $lr=5\cdot 10^{-6}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning. The scene personalization step consists of 1K iterations ($\sim 10$ min.) with a batch size of 1. We set $\lambda=0.1$ in $\mathcal{L}_{loc}$ (Eq. 4).
The novel content personalization step. We extract the foreground object in the reference image using an off-the-shelf segmentation model (Ranftl et al., 2021) and resize it into 512$\times$512. We use the AdamW optimizer with $lr=10^{-4}$, $beta1=0.9$, $beta2=0.999$, $weightdecay=10^{-2}$, and $epsilon=10^{-8}$ for fine-tuning. The scene personalization step consists of 500 iterations ($\sim 5$ min.) with a batch size of 1.
B.3. Details of the coarse editing stage
We implement the coarse editing stage in Pytorch (Paszke et al., 2019). We adopt a sampling strategy of views as in (Zhuang et al., 2023), the resolution of the rendered images is 512$\times$512. We use the Adam optimizer with $beta1=0.9$, $beta2=0.999$, $weightdecay=0$, and $epsilon=10^{-15}$ during optimization. We adopt a linear decay strategy for the learning rate optimizing center position $\mu$ from $10^{-3}$ to $2\cdot 10^{-5}$. As for the other attributes, we use a fixed learning rate the same as the official GS settings (i.e. $10^{-1}$ for opacity $\alpha$, $10^{-2}$ for color $c$, and $5\cdot 10^{-2}$ for the 3D covariance matrix $\Sigma$). Since the complexity of different scenes varies significantly, the optimization iterations are varied from 1K ($\sim 5$ min.) to 5K ($\sim 5$ min.) with batch size of 2. For calculating $\mathcal{L}_{SDS}^{G}$ and $\mathcal{L}_{SDS}^{L}$, we set the CFG weight to 10, and utilize simple two-stage annealing of time step $t$: we sample time steps $t\sim\mathcal{U}(0.02,0.75)$ and anneal into $t\sim\mathcal{U}(0.02,0.25)$.
B.4. Details of the refinement stage
Using results from the coarse editing stage as initialization, we continue optimizing the GS with pixel-level image reconstruction loss, which is applied to the rendered image and the carefully generated pseudo GT image $I_{gt}$. Fig. B.1 illustrates the generation of $I_{gt}$. The optimizer and training hyper-parameters are the same as those in training the initial GS scene (Sec. B.1) except for fewer training iterations (3K).
Appendix C Baselines
Instruct-NeRF2NeRF. Instruct-NeRF2NeRF (“I-N2N”) utilizes Instruct-pix2pix (Brooks et al., 2022) to update the rendered multi-view images according to special text instructions obtained with GPT3 (Brown et al., 2020) for supervision. We implement it using its official code.
We use BLIP-2 (Li et al., 2023) to obtain the textual description of our reference image and then convert it into a suitable instruction as in I-N2N. All the other settings are the same as the default configurations in their official code, including the 3D representation method (Nerfacto (Zhang et al., 2021)), the version of Instruct-pix2pix, and the training hyper-parameters.
DreamEditor. DreamEditor adopts the explicit NeuMesh (Yang et al., 2022) representation distilled from a NeuS (Wang et al., 2021) and includes an attention-based localization operation to support local editing. We implement it using its official code and follow its default training hyperparameters. For fair comparisons, we replace its automatic localization with a more accurate manually selected editing area. Moreover, DreamEditor employs Dreambooth to personalize the target scene. We extend its Dreambooth process to simultaneously personalize both the editing scene and the reference image using two different special tokens in one step. We use the same text prompts as our method to edit the scenes.
Appendix D Evaluation Criteria
We compare our method with the aforementioned baselines on 10 results, covering 3 scenes and different editing types. For each result, we render the edited scene for all methods using the same camera trajectory to obtain video demos, which are used for the user study. All the videos are included in this supplementary.
For metric calculation, we extract 1 image every 5 frames from the above videos and calculate the CLIP-based similarity metric (i.e., CLIP Text-Image directional similarity) and DINO similarity. We report their mean values in Tab. 1 in our main paper.
D.1. CLIP Text-Image directional similarity
CLIP Text-Image directional similarity (Gal et al., 2022b) evaluates the degree of alignment between the change of two images (i.e. before-after editing) and two text prompts (i.e. initial text description $t_{o}$ and the editing text prompt $t_{e}$). Mathematically, it is computed as follows:
(D.1) $\begin{split}\Delta T&=E_{T}(t_{e})-E_{T}(t_{o})\\ \Delta I&=E_{I}(i_{e})-E_{I}(i_{o})\\ CLIP_{dir}&=1-\frac{\Delta I\cdot\Delta T}{|\Delta I||\Delta T|}\end{split}$
where $E_{I}$ and $E_{T}$ are CLIP’s image and text encoder, $i_{o}$ and $i_{e}$ are the original and edited scene image, $t_{o}$ and $t_{e}$ are the texts describing the original and edited scene image. The textual descriptions are obtained with BLIP-2 (i.e. $t_{o}$ for $i_{o}$, $t_{r}$ for the reference image). And we manually generate $t_{e}$ according to $t_{o}$ and $t_{r}$. Note that these text descriptions are not necessarily the same as those used for optimization and do not contain special tokens.
D.2. DINO similarity
Following (He et al., 2023), we use DINO (Oquab et al., 2023) similarity measure the similarity between the edited scene and the reference image from different views.
(D.2) $DINO_{sim}=\frac{E_{D}(i_{e})\cdot E_{D}(i_{r})}{|E_{D}(i_{e})||E_{D}(i_{r})|}$
where $E_{D}$ is the image encoder of DINOv2 (Oquab et al., 2023), and $i_{r}$ the Reference image.

### v3

@@ Abstract.
copyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 121publicationmonth: 7doi: 10.1145/3658205ccs: Computing methodologies Rendering
@@ 2. Related Works
Text-to-image (T2I) diffusion models  (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022), trained on large-scale paired image-text datasets, have gained significant attention since they can generate diverse and high-quality images that match the complicated text prompt. Instead of directly generating images from scratch, another popular and closely related task is to edit the given image according to the text prompt (Meng et al., 2021; Couairon et al., 2022; Kawar et al., 2022; Hertz et al., 2022; Avrahami et al., 2022; Brooks et al., 2022).
Another popular task is object/concept personalization, which aims at generating images for a specified object/concept defined in the given image collection. Textual Inversion (TI) (Gal et al., 2022) optimizes special text token(s) in the text embedding space to represent the specified concept. DreamBooth (Ruiz et al., 2022) fine-tunes the entire diffusion model with a class-specific prior preservation loss as regularization. In general, DreamBooth generates higher-quality images since it involves a larger amount of updated model parameters (i.e. the whole UNet model). However, all the aforementioned methods do not support generating images containing multiple personalized objects simultaneously.
@@
Recently, text-driven radiance field editing methods have gained more and more attention for their editing flexibility and accessibility. For example, Instruct-NeRF2NeRF (Haque et al., 2023) and GenN2N (Liu et al., 2024) employ an image-based diffusion model (InstructPix2Pix (Brooks et al., 2022)) to modify the rendered image by the users’ instructions, and subsequently update the 3D radiance field with the modified image. DreamEditor (Zhuang et al., 2023), 3D Paintbrush (Decatur et al., 2023) and Vox-E (Sella et al., 2023) enable better local editing by adopting explicit 3D representations (i.e. mesh and voxel), where the editing region is automatically determined by the 2D cross-attention maps. GaussianEditor (Chen et al., 2023a; Fang et al., 2023) adopts GS as the scene representation and incorporates 3D semantic segmentation (Kirillov et al., 2023a; Cen et al., 2023) to facilitate efficient and precise scene editing. However, these text-driven approaches lack precise control over the specified appearance and position of the editing results.
@@ 4.1. Stepwise 2D Personalization
In general, our stepwise personalization of the pre-trained T2I model (i.e., Stable Diffusion (SD) (Rombach et al., 2022)) is based on DreamBooth (Ruiz et al., 2022), but with two significant modifications. These changes are essential to personalize both the existing scene and the novel content in the reference image. First, in the 2D personalization of the existing scene, we propose an attention-based localization loss to enforce the interaction between the existing and the novel content specified by the provided 3D bounding box (e.g., sunglasses on the forehead, see Fig. 7). Note that the reference image is not involved in this step. Second, in the 2D personalization of the novel content, we introduce LoRA layers to better capture the unique characteristics of the specified item in the reference image.
@@ 5.3. Comparisons with State-of-the-Art Methods
Quantitative comparisons.
5.4. Ablation Study
Ablation studies on the stepwise 2D personalization
Ablation study on different 3D representations
@@ 6. Conclusion and Limitations
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (NO. 62325605, NO. 62322608), in part by the CAAI-MindSpore Open Fund, developed on OpenI Community,in part by the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (No.VRLAB2023A01).